# Scalable Multi-Task Learning for Particle Collision Event Reconstruction with Heterogeneous Graph Neural Networks

### v1

@@ 4.1 Ablation studies
To quantify the relative merits of the novel architectural developments, several ablation models were trained under similar conditions as described in Section 3.4. Model architectures are either GNN or HGNN and W signifies the use of the weighted message passing introduced in Section 3.2. Ablations further quantify the addition or removal of track node and edge pruning ($\mathcal{L}^{\rm prune}_{\rm BCE}$) and PV association ($\mathcal{L}^{\rm PV}_{\rm BCE}$) tasks by removing the corresponding BCE loss terms from equation 3 during training.
Table 1 shows the various performance metrics including the LCAG loss $\mathcal{L}^{\rm LCA}_{\rm CE}$ and class accuracies as evaluated with the test dataset. Incorporating additional simultaneous tasks consistently improves the accuracy of LCAG reconstruction with the largest gains seen when including track edge and node pruning tasks, which also serve the primary motivation of accelerating inference time as demonstrated in Section 4.2. Although, GNN and HGNN architectures achieve comparable LCAG reconstruction performance, only the HGNN supports PV association, the results of which are presented in Section 4.4. Based on the ablation studies we only present subsequent performance results for GNN and HGNN architectures trained with all possible tasks.
@@ 4.2 Pruning performance
Figure 2(a-b) shows the discrimination power of the track edge and node scores $\hat{y}^{e_{\rm tr}}$ and $\hat{y}^{v_{\rm tr}}$ at HGNN layers 1, 2, 3 and 8. The discrimination power increases significantly between layer 1 and 2 particularly for the edges, with the area under curve (AUC) increasing from 0.973 to 0.999. While the discrimination power saturates with subsequent layers, there are marginal increases in discrimination power with each HGNN layer.
@@
Figure 2: (a) and (b) Receiver Operating Characteristic curves for edge and node scores $\hat{y}^{e_{\rm tr}}$ and $\hat{y}^{v_{\rm tr}}$ at layers 1, 2, 3 and 8. Pruning selections used by subsequent results are highlighted. (c) Relationship between the number of particles from a beauty hadron and the number of particles selected with a tight edge and node pruning $\hat{y}^{e_{\rm tr}}_{8}>0.2$ and $\hat{y}^{v_{\rm tr}}_{8}>0.2$. (d) Signal efficiency and background retention as a function of track multiplicity for tight and loose ($\hat{y}^{e_{\rm tr}}_{8}>0.01$ and $\hat{y}^{v_{\rm tr}}_{8}>0.01$) pruning selections.
Figure 3: GNN and HGNN GPU and CPU mean inference time as a function of track multiplicity per event with and without early pruning ($\hat{y}^{e_{\rm tr}/v_{\rm tr}}_{1-3}>10^{-4}$). The error bands indicate ±$\pm 1\sigma$, where $\sigma$ is the standard deviation of the inference time.
Figure 2(c) and (d) further quantify the performance of the pruning after having applied a selection for the LCAG class pediction to be above zero. In particular, Figure 2(c) displays the relationship between the number of tracks from beauty hadrons and the number of tracks selected with a tight pruning selection, $\hat{y}^{e_{\rm tr}}_{8}>0.2$ and $\hat{y}^{v_{\rm tr}}_{8}>0.2$. Meanwhile, Figure 2(d) compares the efficiencies on signal and background tracks, as a function of track multiplicity for the tight pruning selection and a looser pruning selection with $\hat{y}^{e_{\rm tr}}_{8}>0.01$ and $\hat{y}^{v_{\rm tr}}_{8}>0.01$. The efficiencies are defined by the ratio of surviving signal and background tracks after the pruning and LCAG selections.
Node and edge pruning layers aim to accelerate the inference time of the models, particularly the scaling of inference time with track multiplicity. Figure 3 shows GPU and single-threaded CPU inference times, measured with a NVIDIA RTX 4090 GPU and a Intel Core i9-14900K 3.2 GHz CPU, as a function of track multiplicity for the cases of no pruning selection and a loose edge pruning selection at layers 1-3 ($\hat{y}^{e_{\rm tr}/v_{\rm tr}}_{1-3}>10^{-4}$). Early edge pruning yields noticeable speed-ups above roughly 250 tracks; at multiplicities exceeding 400, we observe a 2-3$\times$ acceleration on GPU and a 5$\times$ acceleration on CPU. For example, at around 300 tracks, single-threaded CPU inference with early pruning takes about 300 ms compared to the 2.2 s timing reported for the multi-stage GNN approach in DFEI [22] on a 2.2 GHz Intel Core processor.  C presents analogous results for batches of eight events.
@@ 4.3 Reconstruction performance
The beauty hadron reconstruction performance is assessed by categorising reconstructed beauty decay chains into several types introduced in  [22]: perfect reconstruction (the reconstructed decay products and the decay hierarchy are both correct and there are no missing decay products, as shown in Figure 1); complete reconstruction (all decay products are correct with none missing; however, the hierarchy is incorrect); not isolated reconstruction (in the reconstruction there are one or more decay products, which are not correct); and partial reconstruction (there are missing decay products). For reconstruction performance we require that the predicted LCAG class is above zero and employ the tight pruning selection, $\hat{y}^{e_{\rm tr}}_{8}>0.2$ and $\hat{y}^{v_{\rm tr}}_{8}>0.2$ introduced in section 4.2.
Table 2 compares the reconstruction performance of the DFEI with the proposed GNN and HGNN models for inclusive beauty hadron collision events. All models significantly outperform the DFEI based on the percentage of perfect hierarchy and complete reconstruction with a much lower occurrence of not isolated decays and a similar level of partial reconstruction. For instance, the HGNN architecture, which slightly outperforms the other models, has a perfect reconstruction 4.8 times that of the DFEI.
Model Perfect reco. Complete reco. Not isolated Part. reco.
@@
Table 2: Comparison of the percentage of each reconstruction category in 10,000 inclusive beauty hadron events using the DFEI method and the GNN/HGNN architectures with tight last layer pruning. Errors show the corresponding statistical uncertainty.
When early pruning is used to accelerate the inference time, there can be a significant drop in performance if pruning cuts are too tight. This is illustrated in Figure 4, which shows the trade-off between reconstruction performance and GPU inference time for track multiplicities above 300 for HGNN and WHGNN models. A gradual tightening of pruning selections in layers 1-3 are applied ranging from $\hat{y}^{e_{\rm tr}}_{1-3}>10^{-7}$ to $\hat{y}^{e_{\rm tr}}_{1-3}>10^{-2}$ in multiples of 10. The WHGNN exhibits a greater invariance of performance with pruning, which allows for smaller inference times at minimal reconstruction performance loss.
Figure 4: Trade off between average GPU timing (ms) and HGNN/WHGNN reconstruction performance for track multiplicity above 300 with no early pruning (the points with the highest GPU timing) and edge pruning in layers 1-3 for selections ranging from $\hat{y}^{e_{\rm tr}}_{1-3}>10^{-7}$ to $\hat{y}^{e_{\rm tr}}_{1-3}>10^{-2}$ in multiples of 10. Shaded histograms of the timing distribution are indicated below.
The reconstruction performance can be further quantified on several exclusive beauty hadron decays with different decay topologies. Table 3 compares the performance of the DFEI model (DF) and the HGNN architecture (H1) on a variety of specific decay modes. While the HGNN often outperforms the DFEI in terms of perfect and complete reconstruction categories and has a smaller fraction of not isolated reconstruction, it suffers from larger levels of partial reconstruction. For comparison, an equivalent HGNN (H2) was trained with the addition of 3000 training samples split uniformly between the decay modes shown in the upper half of Table 3. H2 achieves a much higher perfect hierarchy percentage, while maintaining 1-2% level of partial reconstruction. Moreover, H2 generalises to similar exclusive decays not seen during training, as demonstrated in the lower half of Table 3.
@@
Table 3: Comparison of the beauty hadron reconstruction performance for DFEI and two HGNN models (H1 and H2) with tight pruning, on several exclusive beauty hadron decays. H1 was trained only on inclusive beauty hadron decays while the training of H2 includes a sample of 3000 events split between the decays in the top half the table. We consider the following subdecays $D^{+}[K\pi\pi]$, $D^{-}_{s}[KK\pi]$, $J/\psi[\mu\mu]$, $\phi[KK]$, $K^{*0}[K\pi]$ and $\Lambda^{+}_{c}[pK\pi]$.
An example of the HGNN reconstruction is shown in Figure 5 for an event with seven PVs with one producing $B^{0}\rightarrow D^{+}[K^{+}\pi^{+}\pi^{-}]D^{-}[K^{-}\pi^{+}\pi^{-}]$ and $B^{0}_{s}\rightarrow\bar{p}\pi^{+}\pi^{+}\pi^{-}(\Lambda\rightarrow p\pi^{-})$ decays. Through a combination of integrated pruning and LCAG inference the HGNN is able to isolate the beauty hadron tracks in green and correctly reconstruct the hierarchy of both decay chains.
@@ 4.4 PV association
A key additional capability of the HGNN is PV association, which is indicated in the collision event in Figure 5 by the blue connections between tracks and PVs. This is most valuable for displaced tracks such as those arising from beauty hadron decays and other long-lived particles. Therefore, we quantify the PV-association for all tracks in the event, tracks from the decays of beauty hadrons and the PV-association of whole beauty hadrons. As discussed in section 3.3, a given track is associated to a PV by selecting the PV-track edge for that track with the highest last layer probability score, $\hat{y}^{e_{\rm pv-tr}}_{8}$. Meanwhile, when evaluating the PV-association of the HGNN for beauty hadrons we use the ground-truth to identify the beauty hadron tracks and associate the beauty hadron to the PV to which the majority of its tracks are associated.
@@
HGNN $\mathcal{L}^{\rm PV}_{\rm BCE}$ tr-pv 97.83 ± 0.02 96.19 ± 0.14 97.60 ± 0.15
HGNN $\mathcal{L}^{\rm PV}_{\rm BCE}$ tr-pv, tr-tr 99.83 ± 0.01 99.74 ± 0.04 99.85 ± 0.04
HGNN $\mathcal{L}^{\rm LCA}_{\rm CE},\mathcal{L}^{\rm PV}_{\rm BCE},\mathcal{L}^{\rm prune}_{\rm BCE}$ tr-pv, tr-tr 99.88 ± 0.01 99.78 ± 0.04 99.85 ± 0.04
Table 4: Comparison of the average PV association (%) in inclusive beauty hadron events for various HGNN models with more traditional approaches including association with min IP and a MLP.
Table 4 reports the average PV association accuracy for the conventional “min IP” method, which assigns each track or beauty hadron to the PV with the smallest impact parameter, alongside several HGNN variants, distinguished by their training objectives and edge definitions. Figure 6 illustrates how each method’s track-to-PV association performance varies with the number of reconstructed PVs. Across all scenarios, the HGNN models markedly surpass the min IP baseline, achieving up to 99.8% association accuracy for tracks originating from beauty hadrons when track–track edges are incorporated. A similar figure for beauty hadrons is shown in  D.
Figure 6: Dependence of the average track-to-PV association (%) on PV multiplicity for various methods.
@@ 5 Discussion
The ablation studies in Section 4.1 demonstrate that the inclusion of additional learning objectives noticeably increases the accuracy on the underlying LCAG reconstruction task. This is especially the case for the pruning task as a high discrimination between signal and background nodes is essential for a correct LCAG reconstruction. Having both track edge and node pruning targets at each layer provides important information, helping models learn richer representations that better discriminate between background and signal nodes. Although smaller in impact, the PV association task also improves the LCAG accuracy, since the tracks of a given beauty hadron associate to the same PV.
@@
Another key outcome for practical application is the acceleration in inference time provided by integrated pruning. This enables faster CPU/GPU processing for high track multiplicities with minimal performance loss, which is essential for the model’s scalability as particle collision multiplicity increases in higher luminosity conditions. Here, the proposed weighted message passing scheme, which uses edge and node probability scores, $\hat{y}^{e/v}$, as weights during message passing, allows greater performance invariance when pruning. This can be explained by the learned representations being less dependent on background nodes and edges given their small weighting in message passing.
Although HGNN architectures only deliver a slight gain in beauty hadron reconstruction performance than their GNN counterparts, they provide the extra capability of a precise PV association. The architectural advances proposed in this paper have the potential to significantly enhance PV association performance at LHCb. This improvement would, in turn, lead to greater sensitivity across the LHCb physics program by reducing background contributions and enhancing the resolution of key observables. More accurate PV association would have a particularly strong impact on the precision of measurements involving decays with missing particles, such as neutrinos, by enabling a better determination of the B-hadron flight direction.
@@ 6 Future work
Building on the successful application of HGNNs to beauty hadron reconstruction, several avenues for future development emerge. A major current limitation is the lack of neutral particles, which are potentially challenging to address, primarily because of their lower reconstruction efficiency and resolution compared to charged particles. Within a HGNN, neutral particles can be represented with their own unique representations and learning tasks, which may help overcome the associated difficulties with their inclusion. For instance, regarding multi-task learning, a curriculum learning approach could be employed in which the model first focuses on easier tasks such as learning the LCAG for charged particles and later shifts its attention to tasks involving neutral particles. Beyond incorporating neutrals, several other tasks may be performed to aid reconstruction such as the reconstruction of secondary decay vertices for track-track edges, flavour tagging of reconstructed beauty hadron decays and particle identification of particle nodes.
Several avenues remain for optimisation. First, we did not conduct a comprehensive joint tuning of architectural and training hyperparameters. Second, the edge and node pruning thresholds at each HGNN layer were fixed rather than individually optimised to balance reconstruction accuracy against inference latency. Third, integrating pruning into the training procedure via a “pruning-aware” strategy could yield additional performance gains. Finally, complementary techniques such as model quantisation and weight sparsification could be applied alongside graph pruning to further accelerate inference.
Our current work used a custom simulation environment from DFEI [22], which mimics the Run 3 environment of LHCb. While the simulation considers several reconstruction effects, which are accounted for through experimentally motivated smearing, its limitations (such as the lack of falsely reconstructed tracks and focus on the vertex detector region) necessitate independent confirmation of the promising performance results using official LHCb simulation and eventually data. Furthermore, additional studies are required to investigate the HGNN reconstruction performance for the high luminosity conditions of LHCb’s Upgrade II. Although higher particle and PV multiplicities pose significant challenges, timing information can be incorporated into the heterogenous graph representation to retain performance.

### v2

@@ 4.1 Ablation studies
To quantify the relative merits of the novel architectural developments, several ablation models were trained under similar conditions, as described in Section 3.4. The model architectures are either GNN or HGNN, and W signifies the use of the weighted message passing introduced in Section 3.2. Ablations further quantify the addition or removal of track node and edge pruning ($\mathcal{L}^{\rm prune}_{\rm BCE}$) and PV association ($\mathcal{L}^{\rm PV}_{\rm BCE}$) tasks by removing the corresponding BCE loss terms from Equation 3 during training.
Table 1 shows the various performance metrics, including the LCAG loss $\mathcal{L}^{\rm LCA}_{\rm CE}$ and class accuracies, as evaluated with the test dataset. Incorporating additional tasks consistently improves the accuracy of LCAG reconstruction, with the largest gains seen when including track edge and node pruning tasks. Although the GNN and HGNN architectures achieve comparable LCAG reconstruction performance, only the HGNN supports PV association, the results of which are presented in Section 4.4. Based on the ablation studies we only present subsequent performance results for the GNN and HGNN architectures trained with all possible tasks.
@@ 4.2 Pruning performance
Figure 2(a) and (b) show the discrimination power of the track edge and node scores $\hat{y}^{e_{\rm tr}}$ and $\hat{y}^{v_{\rm tr}}$ at HGNN layers 1, 2, 3 and 8. The discrimination power increases significantly between layers 1 and 2 particularly for the edges, with the area under the curve (AUC) increasing from 0.973 to 0.999. While the discrimination power saturates with subsequent layers, there are marginal increases in the discrimination power for each subsequent HGNN layer.
@@
Figure 2: (a) and (b) Receiver Operating Characteristic curves for edge and node scores $\hat{y}^{e_{\rm tr}}$ and $\hat{y}^{v_{\rm tr}}$ at layers 1, 2, 3 and 8. The pruning selections used in the subsequent results are highlighted. (c) Relationship between the number of particles from a beauty hadron and the number of particles selected with a tight edge and node pruning $\hat{y}^{e_{\rm tr}}_{8}>0.2$ and $\hat{y}^{v_{\rm tr}}_{8}>0.2$. (d) Signal efficiency and background retention as functions of track multiplicity for tight and loose ($\hat{y}^{e_{\rm tr}}_{8}>0.01$ and $\hat{y}^{v_{\rm tr}}_{8}>0.01$) pruning selections.
Figure 3: GNN and HGNN GPU (left) and CPU (right) mean inference time as a function of track multiplicity per event with and without early pruning ($\hat{y}^{e_{\rm tr}/v_{\rm tr}}_{1-3}>10^{-4}$). The error bands indicate ±$\pm 1\sigma$, where $\sigma$ is the standard deviation of the inference time.
Figure 2(c) and (d) further quantify the performance of pruning after applying a selection for the LCAG class pediction to be above zero. In particular, Figure 2(c) displays the relationship between the number of tracks from beauty hadrons and the number of tracks selected with a tight pruning selection, $\hat{y}^{e_{\rm tr}}_{8}>0.2$ and $\hat{y}^{v_{\rm tr}}_{8}>0.2$. Meanwhile, Figure 2(d) compares the efficiencies for signal and background tracks, as a function of track multiplicity using the tight pruning selection and a looser pruning selection with $\hat{y}^{e_{\rm tr}}_{8}>0.01$ and $\hat{y}^{v_{\rm tr}}_{8}>0.01$. The efficiencies are defined by the ratio of surviving signal and background tracks after pruning and LCAG selections.
The node and edge pruning layers aim to accelerate the inference time of the models, particularly the scaling of inference time with track multiplicity. Figure 3 shows GPU and single-threaded CPU inference times, measured with a NVIDIA RTX 4090 GPU and a Intel Core i9-14900K 3.2 GHz CPU, as a function of track multiplicity for the cases of no pruning selection and a loose edge pruning selection at layers 1-3 ($\hat{y}^{e_{\rm tr}/v_{\rm tr}}_{1-3}>10^{-4}$). Early edge pruning yields noticeable speed-ups above approximately 250 tracks; at multiplicities exceeding 400, we observe a 2-3$\times$ acceleration on GPU and a 5$\times$ acceleration on CPU. For example, for events with track multiplicity around 300, single-threaded CPU inference with early pruning takes approximately 300 ms compared to the 2.2 s timing reported for the multi-stage GNN approach in DFEI [22] on a 2.2 GHz Intel Core processor.
@@ 4.3 Reconstruction performance
The beauty-hadron reconstruction performance is evaluated by categorising the reconstructed beauty decay chains into several types introduced in  [22]: perfect reconstruction (the reconstructed decay products and the decay hierarchy are both correct and there are no missing decay products, as shown in Figure 1); complete reconstruction (all decay products are correct with none missing; however, the hierarchy is incorrect); not isolated reconstruction (there are one or more decay products, which are not correct); and partial reconstruction (there are missing decay products). For reconstruction performance, we require that the predicted LCAG class is above zero employ the tight pruning selection, $\hat{y}^{e_{\rm tr}}_{8}>0.2$ and $\hat{y}^{v_{\rm tr}}_{8}>0.2$ introduced in Section 4.2.
Table 2 compares the reconstruction performance of the DFEI with the proposed GNN and HGNN models for inclusive beauty-hadron collision events. All models significantly outperform the DFEI based on the percentage of perfect hierarchy and complete reconstruction with a much lower occurrence of not isolated decays and a similar level of partial reconstruction. For instance, the HGNN architecture, which slightly outperforms the other models, has a perfect reconstruction 4.8 times that of the DFEI.
Model Perfect reco. (%) Complete reco. (%) Not isolated (%) Part. reco. (%)
@@
Table 2: Comparison of the percentage of each reconstruction category in 10,000 inclusive beauty-hadron events using the DFEI method and the GNN/HGNN architectures with tight last layer pruning. Errors show the corresponding statistical uncertainty.
When early pruning is used to accelerate the inference time, there can be a significant drop in performance if the pruning cuts are too tight. This is illustrated in Figure 4, which shows the trade-off between the reconstruction performance and the average GPU inference time for track multiplicities above 300 using the HGNN and WHGNN models. A gradual tightening of pruning selections in layers 1-3 are applied ranging from $\hat{y}^{e_{\rm tr}}_{1-3}>10^{-7}$ to $\hat{y}^{e_{\rm tr}}_{1-3}>10^{-2}$ in multiples of 10. The WHGNN exhibits greater invariance in performance with pruning, which allows for smaller inference times with minimal reconstruction performance loss.
Figure 4: Relationship between the average GPU timing (ms) for track multiplicity above 300 and percentage of each reconstruction category for HGNN (left) and WHGNN (right) models with no early pruning (the points with the highest average GPU timing) and edge pruning in layers 1-3 for selections ranging from $\hat{y}^{e_{\rm tr}}_{1-3}>10^{-7}$ to $\hat{y}^{e_{\rm tr}}_{1-3}>10^{-2}$ in multiples of 10.
The reconstruction performance can be further quantified for several exclusive beauty-hadron decays with different decay topologies. Table 3 compares the performance of the DFEI model (DF) and HGNN architecture (H1) for a variety of specific decay modes. While the HGNN often outperforms the DFEI in terms of perfect and complete reconstruction categories and has a smaller fraction of not isolated reconstruction, it suffers from larger levels of partial reconstruction. For comparison, an equivalent HGNN (H2) was trained with the addition of 3000 training samples split uniformly between the decay modes, which are shown in the upper half of Table 3. H2 achieves a much higher percentage of perfect hierarchy, while maintaining a 1-2% level of partial reconstruction. Moreover, H2 generalises to similar exclusive decays not seen during training, as demonstrated in the lower half of Table 3.
@@
Table 3: Comparison of the percentage of each reconstruction category for DFEI (DF) [22] and two HGNN models (H1 and H2) with tight pruning, for several exclusive beauty-hadron decays. H1 was trained using only inclusive beauty-hadron decays while the training of H2 includes a sample of 3000 events split between the decays in the top half the table. We consider the following subdecays $D^{+}[K\pi\pi]$, $D^{-}_{s}[KK\pi]$, $J/\psi[\mu\mu]$, $\phi[KK]$, $K^{*0}[K\pi]$ and $\Lambda^{+}_{c}[pK\pi]$.
An example of the HGNN reconstruction is shown in Figure 5 for an event with seven PVs with one producing $B^{0}\rightarrow D^{+}[K^{+}\pi^{+}\pi^{-}]D^{-}[K^{-}\pi^{+}\pi^{-}]$ and $B^{0}_{s}\rightarrow\bar{p}\pi^{+}\pi^{+}\pi^{-}[\Lambda\rightarrow p\pi^{-}]$ decays. Through a combination of integrated pruning and LCAG inference, the HGNN can isolate the beauty-hadron tracks in green and correctly reconstruct the hierarchy of both decay chains.
@@ 4.4 PV association
A key additional capability of the HGNN is PV association, which is indicated in the collision event in Figure 5, by the blue connections between tracks and PVs. This is most valuable for displaced tracks, such as those arising from beauty-hadron decays and other long-lived particles. Therefore, we quantified the PV association for all tracks in the event, tracks from the decays of beauty hadrons and the PV-association of whole beauty hadrons. As discussed in Section 3.3, a given track is associated with a PV by selecting the PV-track edge for that track with the highest last-layer probability score, $\hat{y}^{e_{\rm pv-tr}}_{8}$. Meanwhile, when evaluating the PV association of the HGNN for beauty hadrons we use the ground truth to identify the beauty-hadron tracks and associate the beauty hadron with the PV with which the majority of its tracks are associated.
@@
HGNN $\mathcal{L}^{\rm PV}_{\rm BCE}$ pv-tr 97.83 ± 0.02 96.19 ± 0.14 97.60 ± 0.15
HGNN $\mathcal{L}^{\rm PV}_{\rm BCE}$ pv-tr, tr-tr 99.83 ± 0.01 99.74 ± 0.04 99.85 ± 0.04
HGNN $\mathcal{L}^{\rm LCA}_{\rm CE},\mathcal{L}^{\rm PV}_{\rm BCE},\mathcal{L}^{\rm prune}_{\rm BCE}$ pv-tr, tr-tr 99.88 ± 0.01 99.78 ± 0.04 99.85 ± 0.04
Table 4: Comparison of the average PV association (%) in inclusive beauty-hadron events for various HGNN models with the minimum impact parameter method.
Table 4 reports the average PV association accuracy for the conventional “min IP” method, which assigns each track or beauty hadron to the PV with the smallest impact parameter, alongside several HGNN variants, distinguished by their training objectives and edge definitions. Figure 6 illustrates how each method’s PV-track association performance varies with the number of reconstructed PVs. Across all scenarios, the HGNN models markedly surpass the min IP baseline, achieving up to 99.8% association accuracy for tracks originating from beauty hadrons when track–track edges are incorporated. A similar figure for beauty hadrons is shown in  C.
Figure 6: Dependence of the average PV-track association (%) on PV multiplicity for tracks (left) and beauty tracks (right) using various methods.
@@ 5 Discussion
The ablation studies in Section 4.1 demonstrate that the inclusion of additional learning objectives noticeably increases the accuracy of the underlying LCAG reconstruction task. This is especially the case for the pruning task, as a high discrimination between the signal and background nodes is essential for correct LCAG reconstruction. Having both track edge and node pruning targets at each layer provides important information, helping models learn richer representations that better discriminate between background and signal nodes. Although smaller in impact, the PV association task also improves the LCAG accuracy, because the tracks of a given beauty hadron are associated with the same PV.
@@
Another key outcome for practical applications is the acceleration of the inference time provided by integrated pruning. This enables faster CPU/GPU processing for high track multiplicities with minimal performance loss, which is essential for the model’s scalability, as the particle collision multiplicity increases under higher luminosity conditions. The proposed weighted message passing scheme, which uses edge and node probability scores, $\hat{y}^{e/v}$, as weights during message passing, allows greater performance invariance when pruning. This can be explained by the learned representations being less dependent on background nodes and edges given their small weighting in message passing.
Although HGNN architectures only deliver a slight gain in beauty-hadron reconstruction performance compared to their GNN counterparts, they provide the extra capability of precise PV association. The architectural advances proposed in this study have the potential to significantly enhance PV association performance at LHCb. This improvement would, in turn, lead to greater sensitivity across the LHCb physics program by reducing the background contributions and enhancing the resolution of key observables. A more accurate PV association would have a particularly strong impact on the precision of measurements involving decays with missing particles, such as neutrinos, by enabling better determination of the B-hadron flight direction.
@@ 6 Future work
Building on the successful application of HGNNs to beauty hadron reconstruction, several avenues for future development emerge. A major current limitation is the lack of neutral particles, which is potentially challenging to address, primarily because of their lower reconstruction efficiency and resolution compared to charged particles. Within an HGNN, neutral particles can be represented using their own unique representations and learning tasks, which may help overcome the difficulties associated with their inclusion. For instance, in multi-task learning, a curriculum learning approach could be employed in which the model first focuses on easier tasks, such as learning the LCAG for charged particles and later shifts its attention to tasks involving neutral particles. Beyond incorporating neutrals, several other tasks can be performed to aid reconstruction, such as the reconstruction of secondary decay vertices for track-track edges, flavour tagging of reconstructed beauty hadron decays and particle identification of particle nodes.
Several avenues remain for optimisation. First, we did not conduct a comprehensive joint tuning of the architectural and training hyperparameters. Second, the edge and node pruning thresholds at each HGNN layer were fixed rather than individually optimised to balance the reconstruction accuracy against the inference latency. Third, integrating pruning into the training procedure via a “pruning-aware” strategy could yield additional performance gains. Finally, complementary techniques, such as model quantisation and weight sparsification, could be applied alongside graph pruning to further accelerate inference.
In this study we used a custom simulation environment from DFEI [22], which mimics the Run 3 environment of LHCb. While the simulation considers several reconstruction effects, which are accounted for through experimentally motivated smearing, its limitations (such as the lack of falsely reconstructed tracks and focus on the vertex detector region) necessitate independent confirmation of the promising performance results using official LHCb simulation and eventually data. Furthermore, additional studies are required to investigate the HGNN reconstruction performance for the high-luminosity conditions of LHCb’s Upgrade II. Although higher particle and PV multiplicities pose significant challenges, timing information can be incorporated into the heterogeneous graph representation to retain performance.